OpenStudy (anonymous):

four times th sum of -2 and a number is the same as five times the number increased by 1/2

OpenStudy (cruffo):

4 ( n+ (-2) ) = 5n + 1/2

OpenStudy (anonymous):

Find the n value. I am having a hard time working it out. I am getting different values each time.

OpenStudy (cruffo):

Lets distribute the 4 through on the left hand side (and take care of that plus a negative number stuff), \[4n -8 = 5n + \frac{1}{2}\] Then, lets subtract 4n from both sides, \[-8 = n + \frac{1}{2}\] The we can subtract 1/2 from both sides, \[-8 - \frac{1}{2} = n\]

OpenStudy (anonymous):

Great answer thank you!!!

OpenStudy (cruffo):

Make sure to simplify :)